help me please!
a dog trainer has 80 ft of fencing that will be used to create a rectangular work area for dogs . if the trainer wants to enclose an area of 300 ft what will be the dimensions of the work area

Answer by stanbon(75887)

a dog trainer has 80 ft of fencing that will be used to create a rectangular work area for dogs . if the trainer wants to enclose an area of 300 ft^2 what will be the dimensions of the work area
------
Let length be "x"
Then width = (80-2x)/2 = 40-x
-----------------------------------
Equation:
Area = length*width
x(40-x) = 300
-x^2+40x-300 = 0
----
x^2 -40x +300 = 0
----
Factor:
(x-30)(x-10) = 0
-------
x = 10 or x = 30
-----
Length = 10 ft
width = (80-2*10)/2 = 30 ft
===========================
Cheers,
Stan H.